please explain in words and equation.

5. (10 pts) The pH of a 0.350 M solution of uric acid is 2.17. What is the value of Ka?
Answer #1

H+ I HA Uric Acid HA 0.35 + AG o tu ta E 0.35-1 Ka= [H+] [A-] CHA 1² 6035-* • PH = 2017 - PH 10 = 10-2017 [H+] = x = 3) 2=